sammixboo (sammixboo):

hi guys this is austin, sammi's sis and I need help on a ques. Would you make more interest by investing $4,500 in a 6-month certificate of deposit (CD) at 6% interest or by investing $3,500 in a 12-month CD at 4% interest?

OpenStudy (anonymous):

Without more information I would assume this is simple interest. I = Prt P is the initial investment (4,500 or 3,500) r is the annual interest rate (6% or 4%) t is number of years (1/2 or 1 for 6 and 12 months) So solve the two equations. Whichever has a higher solution is your answer.

OpenStudy (anonymous):

I missed that this is asking which has a higher interest gained. When you solve the Prt equation, then subtract what was put in, that would tell you what the interest alone was.

sammixboo (sammixboo):

ok I will and I will show you what I did and can you tell me if I did it right?

OpenStudy (anonymous):

Sure.

sammixboo (sammixboo):

this is what I put First section: I = prt I = 4,500*6%*1/2 I = 4,500*0.06*0.5 I = 270*1/2 I = 270*0.5 I = 135 ---------------------- Second section: I = prt I = 3,500*4%*1 I = 3,500*0.04*1 I = 140*1 I = 140 ---------------------------- So if you invest $3,500 in a 12-month CD at 4% interest is will have more interest than investing $4,500 in a 6-month (CD) at 6%.

OpenStudy (anonymous):

With the information given, I would say yes, you are right.

sammixboo (sammixboo):

Thanks and yes he got it right.

random231 (random231):

haha silly